OpenStudy (mathmath333):

Spiderman crawls 12 m up and falls back 5 m every alternate hour.if the building is 63 m high, how long will it take spiderman to reach the top.

OpenStudy (welshfella):

so first 1hour he climbs up 7 m second hour he climbs 12 m and doesn't fall back right?

OpenStudy (mathmath333):

well its kind of confusing

OpenStudy (welshfella):

yea it is maybe he climbs 12 m first hour then 7 m the second!

OpenStudy (mathmath333):

u may assume as u think is right.

OpenStudy (welshfella):

ok then sequence is 7 + 12 + 7 + 12 + 7 + 12 that's 57 m in 6 hours so he has 6 m to go he can climb this in 0.5 hours so i guess that is 6.5 hours

OpenStudy (welshfella):

if we take the other scenario 12+7+12+7+12+ 7 = 57 its the same right?

OpenStudy (mathmath333):

lest me give u options a.) 18 hrs b.) 17 hrs c.) 16 hrs 30 min d.) 16 hrs 40 minutes

OpenStudy (mathmath333):

ur "procedure" is correct

OpenStudy (welshfella):

maybe he doesn't do the last 6 meters in one go - I dont know to be honest and I have to go

OpenStudy (mathmath333):

ok , i think it just depends on the interpretation of the question.

ganeshie8 (ganeshie8):

18 hrs ?

OpenStudy (mathmath333):

no, haha

ganeshie8 (ganeshie8):

it better be 17 hrs then

OpenStudy (mathmath333):

nono

OpenStudy (mathmath333):

lol are u trying every option

ganeshie8 (ganeshie8):

12*9 -5*8 = 68 this gives me 9 ups (9 hours) and 8 downs (8 hours) 9 + 8 = 17 hrs

OpenStudy (mathmath333):

look at the question its 63 not 68

ganeshie8 (ganeshie8):

12*8 - 5*8 = 56 is too short

OpenStudy (mathmath333):

well i gtg now , this is a nice little tricky question.

ganeshie8 (ganeshie8):

it maybe a nice question, but the question has to be phrased unambiguously

OpenStudy (mathmath333):

yes but options make it clearer i think.

ganeshie8 (ganeshie8):

it could be anything based on how we interpret the problem

ganeshie8 (ganeshie8):

i gave you justification for 17hrs earlier

ganeshie8 (ganeshie8):

if the minimum time interval is assumed to be 5 minutes, we get 16hrs 35 mnts which is not even there in options :/

OpenStudy (mathmath333):

yes but 12*9 -5*8 = \(68\) but the building is \(63\) m high.

ganeshie8 (ganeshie8):

17hrs is right if we assume the minimum time interval is 1 hour

OpenStudy (mathmath333):

lol actually 16 hrs and 35 min is correct haha

ganeshie8 (ganeshie8):

the question doesn't clearly specify whether the spiderman crawls continuously or he leaps 12 m every one hour.

OpenStudy (mathmath333):

thats what makes this question tricky

ganeshie8 (ganeshie8):

Nope. tricky questions are fun, poorly phrased questions are not :P

OpenStudy (mathmath333):

otherwise it would be no fun. i also failed first time

OpenStudy (mathmath333):

it may be poorly phrased but it is funny.